The atomic density of a solid is 5.85 ×1028 m−3. Its electrical resistivity is 1.6 ×10−8Ωm. Assume that each atom contributes one conduction electron:
The drift mobility (in m2 V−1 s−1) of the conduction electrons is?
1
Expert's answer
2021-09-19T09:25:26-0400
"\mu=\frac{1}{ne\rho}\\\mu=\frac{1}{(1.6\cdot10^{-8})(5.85\cdot10^{28})(1.6\cdot10^{-19})}=0.00668\frac{m^2}{VS}"